Can you please help me with both questions please
4 -2.16 points BBBasicSta:7 3 R.004. Radon: The Problem No One Wants to Face is the title of an article appearing in Consumer Reports. Radon is a gas emitted from the ground that can collect in houses and buildings. At certain levels it can cause lung cancer. Radon concentrations are measured in picocuries per litr (pC/L). A radon level of 4 pCV/L is considered acceptable. Radon levels in a house vary from week to week. In one house, a sample of 8 weeks had the following readings for radon level (in pCyL). 1.9 2.7 5.7 4.2 1.9 8.1 3.9 7.5 (e) Find the mean, median, and mode. (Round your answers to two decimal places.) mean 4.49 median 4.05 mode 1.90 (b) Find the sample standard deviation, coefficient of variation, and range. (Round your answers to two decimal places.) cV range
media%2Fedf%2Fedf628da-984a-491f-b38d-9c
media%2F825%2F8256e4a9-da85-4b3f-8175-94
Answer #1

Answer 4

Calculation for Standard Deviation is as follows:

8.1 65.61 56.25 32.49 17.64 15.21 7.29 3.61 3.61 201.71 5.7 4.2 3.9 2.7 35.9 The sample mean X is computed as follows 35.9 Ai

Therefore, s = 2.41

Coefficient of Variation = Standard Deviation/Mean = 2.41/4.49 = 0.5367 = 53.67%

Range = Largest Value - Smallest Value = 8.1 - 1.9 = 6.2

Answer 5

(a) Correct option is "Stratified Sample"

Reason: Stratified sample is a sample that is drawn from a number of separate strata of the population, rather than at random from the whole population, in order that it should be representative.

(b) Correct option is "Student on your campus with work study jobs"

It is given in question directly

(c) Correct option is "Hours scheduled"

It is given in question directly

(d) Correct option is "Both Quantitative and Qualitative"

Quatitative: Hours scheduled

Qualitative: Applicability of work experience to future employment goals

We could only answer all parts of Question 1 and 4 subparts of Question 2 in given time period